# Beamfocusing Optimization for Near-Field Wideband Multi-User Communications

### v4

@@ III-B2 Subproblem With Respect to $\mathbf{D}_{m}$
The optimal solution of $\mathbf{D}_{m}$ can be obtained by minimizing the second term of the objective function in problem (III-A) without constraint, which is given by
@@ III-B3 Subproblem With Respect to $\mathbf{W}_{m}$
$\max_{\mathbf{W}_{m}}\quad$ $\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}\log_{2}(1+\hat{\gamma}_{m,k}(\mathbf{W}_{m}))$
$-\frac{1}{\rho}\sum_{m\in\mathcal{M}}\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2}.$ (33)
While this problem also has no constraint, its non-convex objective function makes it still challenging to solve. To address this issue, we adopt the quadratic transform [28, Theorem 2] and Lagrangian dual transform [29, Theorem 3] to convert problem (III-B3) into the following equivalent form:
$\max_{\mathbf{W}_{m},\bm{\lambda}_{m},\bm{\mu}_{m}}$ $\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}\Bigg{(}2\sqrt{1+\mu_{m,k}}\mathrm{Re}\left\{\lambda_{m,k}^{*}\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,k}\right\}$
$-|\lambda_{m,k}|^{2}\left(\sum_{i\in\mathcal{K}}|\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,i}|^{2}+\frac{\sigma_{m,k}^{2}}{P_{t}}\|\mathbf{W}_{m}\|_{F}^{2}\right)\Bigg{)}$
$-\frac{1}{\rho}\sum_{m\in\mathcal{M}}\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2}\triangleq g(\mathbf{W}_{m},\bm{\lambda}_{m},\bm{\mu}_{m}),$ (34)
where $\bm{\lambda}_{m}=[\lambda_{m,1},\lambda_{m,2},\dots,\lambda_{m,K}]^{T}$ and $\bm{\mu}_{m}=[\mu_{m,1},\mu_{m,2},\dots,\mu_{m,K}]^{T}$ are auxiliary variables. Following [28] and [29], for any given $\mathbf{W}_{m}$, $\mu_{m,k}$ and $\lambda_{m,k}$ can be sequentially updated by the following closed-form solution:
@@
$\lambda_{m,k}=$ $\frac{\sqrt{1+\mu_{m,k}}\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,k}}{\sum_{i\in\mathcal{K}}|\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,i}|^{2}+\frac{\sigma_{m,k}^{2}}{P_{t}}\|\mathbf{W}_{m}\|_{F}^{2}},\forall m,k.$ (36)
Then, for any given $\bm{\lambda}_{m}$ and $\bm{\mu}_{m}$, problem (III-B3) is an unconstrained convex optimization problem for $\mathbf{W}_{m}$. To obtain the optimal solution, we first derive the first-order derivative of $g(\mathbf{W}_{m},\bm{\lambda}_{m},\bm{\mu}_{m})$ with respect to $\mathbf{W}_{m}$ as follows:
$\frac{\partial g(\mathbf{W}_{m},\bm{\lambda}_{m},\bm{\mu}_{m})}{\partial\mathbf{W}_{m}}=\mathbf{W}_{m}^{H}\Bigg{(}\frac{1}{\rho}\mathbf{I}_{N}+\sum_{k\in\mathcal{K}}|\lambda_{m,k}|^{2}\Big{(}\mathbf{h}_{m,k}\mathbf{h}_{m,k}^{H}$
$+\frac{\sigma_{m,k}^{2}}{P_{t}}\mathbf{I}_{N}\Big{)}\Bigg{)}-\frac{1}{\rho}\mathbf{D}_{m}^{H}\mathbf{T}_{m}^{H}\mathbf{A}^{H}-\sum_{k\in\mathcal{K}}\sqrt{1+\mu_{m,k}}\lambda_{m,k}\mathbf{e}_{k}\mathbf{h}_{m,k}^{H},$ (37)
where $\mathbf{e}_{k}\in\mathbb{R}^{K\times 1}$ is a vector whose $k$-th entry is one while the others are zero. Based on the first-order optimality condition $\frac{\partial g(\mathbf{W}_{m},\bm{\lambda}_{m},\bm{\mu}_{m})}{\partial\mathbf{W}_{m}}=0$, the optimal $\mathbf{W}_{m}$ is given by
$\mathbf{W}_{m}=\Bigg{(}\frac{1}{\rho}\mathbf{I}_{N}+\sum_{k\in\mathcal{K}}|\lambda_{m,k}|^{2}\Big{(}\mathbf{h}_{m,k}\mathbf{h}_{m,k}^{H}+\frac{\sigma_{m,k}^{2}}{P_{t}}\mathbf{I}_{N}\Big{)}\Bigg{)}^{-1}$
×$\times\left(\frac{1}{\rho}\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}+\sum_{k\in\mathcal{K}}\sqrt{1+\mu_{m,k}}\lambda_{m,k}^{*}\mathbf{h}_{m,k}\mathbf{e}_{k}^{T}\right).$ (38)
@@ III-B4 Overall Algorithm, Convergence, and Complexity
Building upon the solutions developed for each optimization block, the principles of the penalty method, and the results in Lemma 1, the overall penalty-based FDA algorithm for solving problem (II-C) for fully-connected architectures is summarized in Algorithm 2. Since the objective value is non-increasing in each step of block coordinate descent in the inner loop and is bounded from below, the convergence of Algorithm 2 is guaranteed. Furthermore, Algorithm 2 is computationally efficient, as the optimization variables in each step are updated by either the closed-form solution or the low-complexity one-dimensional search. It can be shown that the complexities of updating $\mathbf{A}$, $\mathbf{T}_{m}$, and $\mathbf{V}_{m}$ in Algorithm 1 are in order of $\mathcal{O}(MN)$, $\mathcal{O}(MN^{2}N_{\mathrm{RF}}Q/N_{\mathrm{T}})$, and $\mathcal{O}(M(NKN_{\mathrm{RF}}+(NN_{\mathrm{T}}+K)N_{\mathrm{RF}}^{2}+N_{\mathrm{RF}}^{3}))$, respectively, where $\mathcal{O}(\cdot)$ is the big-O notation. Furthermore, the complexities for updating $\mathbf{D}_{m}$, $\bm{\mu}_{m,k}$, $\bm{\lambda}_{m,k}$, and $\mathbf{W}_{m}$ in Algorithm 2 are in order of $\mathcal{O}(M(2NN_{\mathrm{RF}}^{2}+N_{\mathrm{RF}}^{3}+NN_{\mathrm{RF}}K))$, $\mathcal{O}(MK(N+K(K-1)N+N^{2}))$, $\mathcal{O}(MK((K^{2}+1)N+N^{2}))$, and $\mathcal{O}(M((KN_{\mathrm{RF}}+K^{2})N+KN^{2}+N^{3}))$, respectively.
@@
8:     until the fractional increase of the objective value of problem (III-A) falls below a predefined threshold
@@
10:  until the penalty value falls below a predefined threshold
11:  scale the digital beamformer as $\mathbf{D}_{m}=\frac{\sqrt{P_{t}}}{\|\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2}}\tilde{\mathbf{D}}_{m}$
@@ III-C Solution for Sub-Connected Architectures
For the sub-connected architecture, the BCD method can also be employed to solve problem (III-A). It can be shown that for fixed $\mathbf{A}$ and $\mathbf{T}_{m}$, the blocks $\mathbf{D}_{m}$ and $\mathbf{W}_{m}$ can be updated using the solutions in (III-B2) and (III-B3), respectively. Therefore, in this subsection, we focus on the subproblem with respect to $\mathbf{A}$ and $\mathbf{T}_{m}$, which is given by
$\min_{\mathbf{A},\mathbf{T}_{m}}\quad$ $\sum_{m\in\mathcal{M}}\left\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\right\|_{F}^{2}$ (39a)
@@
Unlike the fully-connected architecture, the analog beamforming matrix $\mathbf{A}\mathbf{T}_{m}$ in the sub-connected architecture is a block-diagonal matrix rather than a full matrix according to (11) and (12). By exploiting this spectral structure of matrix $\mathbf{A}\mathbf{T}_{m}$, problem (III-C) can be directly solved without introducing additional auxiliary variable $\mathbf{V}_{m}$. In particular, according to (11), the objective function can be equivalently written as
(39a)$=\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{m\in\mathcal{M}}\left\|\mathbf{\Psi}_{m,n}-\mathbf{A}_{n}e^{-j2\pi f_{m}\mathbf{t}_{n}}\mathbf{p}_{m,n}^{H}\right\|_{F}^{2}$
$=\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{m\in\mathcal{M}}\Big{(}\mathbf{p}_{m,n}^{H}\mathbf{p}_{m,n}\left(e^{-j2\pi f_{m}\mathbf{t}_{n}}\right)^{H}\mathbf{A}_{n}^{H}\mathbf{A}_{n}e^{-j2\pi f_{m}\mathbf{t}_{n}}$
$\hskip 14.22636pt+\mathrm{tr}\left(\mathbf{\Psi}_{m,n}\mathbf{\Psi}_{m,n}^{H}\right)-\mathrm{Re}\left\{\mathbf{p}_{m,n}^{H}\mathbf{\Psi}_{m,n}^{H}\mathbf{A}_{n}e^{-j2\pi f_{m}\mathbf{t}_{n}}\right\}\Big{)},$ (40)
where $\mathbf{\Psi}_{m,n}=\mathbf{W}_{m}((n-1)N_{\mathrm{sub}}+1:nN_{\mathrm{sub}},:)$ and $\mathbf{p}_{m,n}=\mathbf{D}_{m}^{H}(:,n)$. Based on (12), we further have
$\left(e^{-j2\pi f_{m}\mathbf{t}_{n}}\right)^{H}\mathbf{A}_{n}^{H}\mathbf{A}_{n}e^{-j2\pi f_{m}\mathbf{t}_{n}}$
$=\sum_{l=1}^{N_{\mathrm{T}}}e^{j2\pi f_{m}t_{n,l}}\mathbf{a}_{n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}=N_{\mathrm{sub}},$ (41)
where the last step stems from the unit-modulus constraint on $\mathbf{a}_{n,l}$. Based on the above result, it can be concluded that only the last term in (III-C) is related to the optimization variable of interest, which can be further simplified as
$\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{m\in\mathcal{M}}\mathrm{Re}\left\{\mathbf{p}_{m,n}^{H}\mathbf{\Psi}_{m,n}^{H}\mathbf{A}_{n}e^{-j2\pi f_{m}\mathbf{t}_{n}}\right\}$
$=$ $\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{l=1}^{N_{\mathrm{T}}}\sum_{m\in\mathcal{M}}\mathrm{Re}\left\{\bm{\phi}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\},$ (42)
where $\bm{\phi}_{m,n,l}=\mathbf{\Phi}_{m,n}\left((l-1)\frac{N_{\mathrm{sub}}}{N_{\mathrm{T}}}+1:l\frac{N_{\mathrm{sub}}}{N_{\mathrm{T}}}\right)$ and $\mathbf{\Phi}_{m,n}=\mathbf{\Psi}_{m,n}\mathbf{p}_{m,n}$.
$\max_{\mathbf{a}_{n,l},t_{n,l}}\quad$ $\sum_{m\in\mathcal{M}}\mathrm{Re}\left\{\bm{\phi}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}\triangleq\tilde{h}(\mathbf{a}_{n,l},t_{n,l})$ (43a)
$\mathrm{s.t.}\quad$ $|\mathbf{a}_{n,l}(i)|=1,\forall i,$ (43b)
$t_{n,l}\in[0,t_{\max}].$ (43c)
For any given $t_{n,l}$, the optimal solution of $\mathbf{a}_{n,l}$ is given by
$\mathbf{a}_{n,l}=e^{j\angle\left(\sum_{m\in\mathcal{M}}\bm{\phi}_{m,n,l}e^{j2\pi f_{m}t_{n,l}}\right)}.$ (44)
$t_{n,l}=\arg\max_{t_{n,l}\in\mathcal{Q}}\tilde{h}(\mathbf{a}_{n,l},t_{n,l}).$ (45)
@@
9:     until the fractional increase of the objective value of problem (III-A) falls below a predefined threshold
@@
11:  until the penalty value falls below a predefined threshold
12:  scale the digital beamformer as $\mathbf{D}_{m}=\frac{\sqrt{P_{t}}}{\|\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2}}\tilde{\mathbf{D}}_{m}$
@@
The FDA approach based on BCD for sub-connected architectures is summarized in Algorithm 3. The convergence of this algorithm is guaranteed because the objective value is non-increasing in each step of block coordinate descent in the inner loop and is bounded from below. In each iteration, the complexities of updating $\mathbf{A}$ and $\mathbf{T}_{m}$ are given by $\mathcal{O}(M(N^{2}K^{2}+N))$ and $\mathcal{O}(MN^{2}Q/(N_{\mathrm{RF}}N_{\mathrm{T}}))$, respectively, while the complexities of updating $\mathbf{D}_{m}$, $\bm{\mu}_{m,k}$, $\bm{\lambda}_{m,k}$, and $\mathbf{W}_{m}$ are the same as the counterpart in Algorithm 2.
IV Low-complexity Heuristic Two-stage Approach

### v6

@@ III-B2 Subproblem With Respect to $\mathbf{D}_{m}$
The optimal solution of $\mathbf{D}_{m}$ can be obtained by minimizing the second term of the objective function in problem (21) without constraint, which is given by
@@ III-B3 Subproblem With Respect to $\mathbf{W}_{m}$
$\max_{\mathbf{W}_{m}}\quad$ $\sum_{m=1}^{M}\sum_{k=1}^{K}\log_{2}(1+\hat{\gamma}_{m,k}(\mathbf{W}_{m}))$
$-\frac{1}{\rho}\sum_{m=1}^{M}\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2}.$ (33)
While this problem also has no constraint, its non-convex objective function makes it still challenging to solve. To address this issue, we adopt the quadratic transform [33, Theorem 2] and Lagrangian dual transform [34, Theorem 3] to convert problem (III-B3) into the following equivalent form:
$\max_{\mathbf{W}_{m},\bm{\lambda}_{m},\bm{\mu}_{m}}\sum_{m=1}^{M}\sum_{k=1}^{K}\Bigg{(}2\sqrt{1+\mu_{m,k}}\Re\left\{\lambda_{m,k}^{*}\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,k}\right\}$
$-|\lambda_{m,k}|^{2}\left(\sum_{i=1}^{K}|\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,i}|^{2}+\frac{\sigma_{m,k}^{2}}{P_{t}}\|\mathbf{W}_{m}\|_{F}^{2}\right)\Bigg{)}$
$-\frac{1}{\rho}\sum_{m=1}^{M}\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2}\triangleq g(\mathbf{W}_{m},\bm{\lambda}_{m},\bm{\mu}_{m}),$ (34)
where $\bm{\lambda}_{m}=[\lambda_{m,1},\lambda_{m,2},\dots,\lambda_{m,K}]^{T}$ and $\bm{\mu}_{m}=[\mu_{m,1},\mu_{m,2},\dots,\mu_{m,K}]^{T}$ are auxiliary variables. Following [33] and [34], for any given $\mathbf{W}_{m}$, the values of $\mu_{m,k}$ and $\lambda_{m,k}$ can be sequentially updated by the following closed-form solution:
@@
$\lambda_{m,k}=$ $\frac{\sqrt{1+\mu_{m,k}}\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,k}}{\sum_{i=1}^{K}|\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,i}|^{2}+\frac{\sigma_{m,k}^{2}}{P_{t}}\|\mathbf{W}_{m}\|_{F}^{2}},\forall m,k.$ (36)
Then, for any given $\bm{\lambda}_{m}$ and $\bm{\mu}_{m}$, problem (III-B3) is an unconstrained convex optimization problem for $\mathbf{W}_{m}$. Based on the first-order optimality condition $\frac{\partial g(\mathbf{W}_{m},\bm{\lambda}_{m},\bm{\mu}_{m})}{\partial\mathbf{W}_{m}}=0$, the optimal $\mathbf{W}_{m}$ is given by
$\mathbf{W}_{m}=\Bigg{(}\frac{1}{\rho}\mathbf{I}_{N}+\sum_{k=1}^{K}|\lambda_{m,k}|^{2}\Big{(}\mathbf{h}_{m,k}\mathbf{h}_{m,k}^{H}+\frac{\sigma_{m,k}^{2}}{P_{t}}\mathbf{I}_{N}\Big{)}\Bigg{)}^{-1}$
×$\times\left(\frac{1}{\rho}\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}+\sum_{k=1}^{K}\sqrt{1+\mu_{m,k}}\lambda_{m,k}^{*}\mathbf{h}_{m,k}\mathbf{e}_{k}^{T}\right),$ (37)
where $\mathbf{e}_{k}\in\mathbb{R}^{K\times 1}$ is a vector whose $k$-th entry is one while the others are zero.
@@ III-B4 Overall Algorithm, Convergence, and Complexity
Building upon the solutions developed for each optimization block, the principles of the penalty method, and the results in Lemma 1, the overall penalty-based FDA algorithm for solving problem (15) for fully-connected architectures is summarized in Algorithm 2. Since the objective value is non-increasing in each step of block coordinate descent in the inner loop and is bounded from below, the convergence of Algorithm 2 is guaranteed. Furthermore, Algorithm 2 is computationally efficient, as the optimization variables in each step are updated by either the closed-form solution or the low-complexity one-dimensional search. It can be shown that the complexities of updating $\mathbf{A}$, $\mathbf{T}_{m}$, and $\mathbf{V}_{m}$ in Algorithm 1 are in order of $\mathcal{O}(MN)$, $\mathcal{O}(MN^{2}N_{\mathrm{RF}}Q/N_{\mathrm{T}})$, and $\mathcal{O}(M(NKN_{\mathrm{RF}}+(NN_{\mathrm{T}}+K)N_{\mathrm{RF}}^{2}+N_{\mathrm{RF}}^{3}))$, respectively, where $\mathcal{O}(\cdot)$ is the big-O notation. Furthermore, the complexities for updating $\mathbf{D}_{m}$, $\bm{\mu}_{m,k}$, $\bm{\lambda}_{m,k}$, and $\mathbf{W}_{m}$ in Algorithm 2 are in order of $\mathcal{O}(M(2NN_{\mathrm{RF}}^{2}+N_{\mathrm{RF}}^{3}+NN_{\mathrm{RF}}K))$, $\mathcal{O}(MK(N+K(K-1)N+N^{2}))$, $\mathcal{O}(MK((K^{2}+1)N+N^{2}))$, and $\mathcal{O}(M((KN_{\mathrm{RF}}+K^{2})N+KN^{2}+N^{3}))$, respectively.
@@
8:     until the fractional increase of the objective value of problem (21) falls below a threshold
@@
10:  until the penalty value falls below a threshold
11:  scale the digital beamformer as $\mathbf{D}_{m}=\frac{\sqrt{P_{t}}}{\|\tilde{\mathbf{A}}\tilde{\mathbf{T}}_{m}\tilde{\mathbf{D}}_{m}\|_{F}^{2}}\tilde{\mathbf{D}}_{m}$
@@ III-C Solution for Sub-connected Architectures
For the sub-connected architecture, the BCD method can also be employed to solve problem (21). It can be shown that for fixed $\mathbf{A}$ and $\mathbf{T}_{m}$, the blocks $\mathbf{D}_{m}$ and $\mathbf{W}_{m}$ can be updated using the solutions in (III-B2) and (III-B3), respectively. Therefore, in this subsection, we focus on the subproblem with respect to $\mathbf{A}$ and $\mathbf{T}_{m}$, which is given by
$\min_{\mathbf{A},\mathbf{T}_{m}}\quad$ $\sum_{m=1}^{M}\left\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\right\|_{F}^{2}$ (38a)
@@
Unlike the fully-connected architecture, the analog beamforming matrix $\mathbf{A}\mathbf{T}_{m}$ in the sub-connected architecture is a block-diagonal matrix rather than a full matrix according to (11) and (12). By exploiting this spectral structure of matrix $\mathbf{A}\mathbf{T}_{m}$, problem (38) can be directly solved without introducing additional auxiliary variable $\mathbf{V}_{m}$. In particular, according to (11), the objective function can be equivalently written as
(38a)$=\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{m=1}^{M}\left\|\mathbf{\Psi}_{m,n}-\mathbf{A}_{n}e^{-j2\pi f_{m}\mathbf{t}_{n}}\mathbf{p}_{m,n}^{H}\right\|_{F}^{2}$
$=\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{m=1}^{M}\Big{(}\Xi_{m,n}\mathbf{p}_{m,n}^{H}\mathbf{p}_{m,n}+\mathrm{tr}\left(\mathbf{\Psi}_{m,n}\mathbf{\Psi}_{m,n}^{H}\right)$
$\hskip 51.21504pt-\Re\left\{\mathbf{p}_{m,n}^{H}\mathbf{\Psi}_{m,n}^{H}\mathbf{A}_{n}e^{-j2\pi f_{m}\mathbf{t}_{n}}\right\}\Big{)},$ (39)
where $\mathbf{\Psi}_{m,n}=\mathbf{W}_{m}((n-1)N_{\mathrm{sub}}+1:nN_{\mathrm{sub}},:)$, $\mathbf{p}_{m,n}=\mathbf{D}_{m}^{H}(:,n)$, and
$\Xi_{m,n}=\sum_{l=1}^{N_{\mathrm{T}}}\mathbf{a}_{n,l}^{H}\mathbf{a}_{n,l}e^{j2\pi f_{m}t_{n,l}}e^{-j2\pi f_{m}t_{n,l}}=N_{\mathrm{sub}},$ (40)
where is obtained based on (12) and the unit-modulus constraint on $\mathbf{a}_{n,l}$. Based on the above result, it can be concluded that only the last term in (III-C) is related to the optimization variable of interest, which can be further simplified as
$\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{m=1}^{M}\Re\left\{\mathbf{p}_{m,n}^{H}\mathbf{\Psi}_{m,n}^{H}\mathbf{A}_{n}e^{-j2\pi f_{m}\mathbf{t}_{n}}\right\}$
$=$ $\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{l=1}^{N_{\mathrm{T}}}\sum_{m=1}^{M}\Re\left\{\bm{\phi}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\},$ (41)
where $\bm{\phi}_{m,n,l}=\mathbf{\Phi}_{m,n}\big{(}(l-1)\frac{N_{\mathrm{sub}}}{N_{\mathrm{T}}}+1:l\frac{N_{\mathrm{sub}}}{N_{\mathrm{T}}}\big{)}$ and $\mathbf{\Phi}_{m,n}=\mathbf{\Psi}_{m,n}\mathbf{p}_{m,n}$.
@@
9:     until the fractional increase of the objective value of problem (21) falls below a threshold
@@
11:  until the penalty value falls below a threshold
12:  scale the digital beamformer as $\mathbf{D}_{m}=\frac{\sqrt{P_{t}}}{\|\tilde{\mathbf{A}}\tilde{\mathbf{T}}_{m}\tilde{\mathbf{D}}_{m}\|_{F}^{2}}\tilde{\mathbf{D}}_{m}$
$\max_{\mathbf{a}_{n,l},t_{n,l}}\quad$ $\sum_{m=1}^{M}\Re\left\{\bm{\phi}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}$ (42a)
$\mathrm{s.t.}\quad$ $|\mathbf{a}_{n,l}(i)|=1,\forall i,$ (42b)
$t_{n,l}\in[0,t_{\max}].$ (42c)
For any given $t_{n,l}$, the optimal solution of $\mathbf{a}_{n,l}$ is given by
$\mathbf{a}_{n,l}=e^{j\angle\left(\sum_{m=1}^{M}\bm{\phi}_{m,n,l}e^{j2\pi f_{m}t_{n,l}}\right)}.$ (43)
$t_{n,l}=\arg\max_{t_{n,l}\in\mathcal{Q}}\sum_{m=1}^{M}\Re\left\{\bm{\phi}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}.$ (44)
@@
The FDA approach based on BCD for sub-connected architectures is summarized in Algorithm 3. The convergence of this algorithm is guaranteed because the objective value is non-increasing in each step of block coordinate descent in the inner loop and is bounded from below. In each iteration, the complexities of updating $\mathbf{A}$ and $\mathbf{T}_{m}$ are given by $\mathcal{O}(M(N^{2}K^{2}+N))$ and $\mathcal{O}(MN^{2}Q/(N_{\mathrm{RF}}N_{\mathrm{T}}))$, respectively, while the complexities of updating $\mathbf{D}_{m}$, $\bm{\mu}_{m,k}$, $\bm{\lambda}_{m,k}$, and $\mathbf{W}_{m}$ are the same as the counterpart in Algorithm 2. We note that the main complexity of the proposed Algorithm 2 and Algorithm 3 stems from the matrix multiplication and inversion for calculating the closed-form updates over a large number of iterations. To accelerate the algorithm and facilitate its practical implementation, deep unfolding can be employed to transform the iterative algorithm into a trainable deep neural network with a small number of layers, based on the structure of the closed-form updates in each iteration [35, 36].
IV Heuristic Two-stage Approach